# Polarity transitions induced by symmetry-breaking outer boundary heat flux in rapidly rotating dynamos

### v1

@@ Abstract
This study investigates, analytically and numerically, the role of an equatorially anti-symmetric lateral variation in heat flux at the outer boundary in polarity transitions in rapidly rotating dynamos. In an unstably stratified fluid, the frequencies of vertical and horizontal (lateral) buoyancy complement each other such that a polarity transition is induced by the boundary anti-symmetry through the suppression of the slow magnetic-Archimedean-Coriolis (MAC) waves at relatively small vertical buoyant forcing. A dipole-dominated dynamo in the low-inertia limit transitions to polarity-reversing and multipolar states in succession as the relative intensity of horizontal buoyancy is progressively increased for a fixed vertical buoyancy. In the same parameter space, an equatorially symmetric heat flux variation does not induce a polarity transition. A composite boundary heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a horizontal buoyancy of the same order as that for the transition induced by a purely anti-symmetric variation. This makes the analysis relevant to Earth’s core, which convects in response to a composite heat flux variation in the lowermost mantle. A heterogeneity with a dominant symmetric part inhibits polarity transitions, likely producing long periods in Earth’s history without reversals. The fact that compositional buoyancy is much stronger than thermal buoyancy, together with the known order of magnitude of the peak field intensity in the inertia-free limit, indicates that the core can operate under a large lower-mantle heat flux heterogeneity, of $O(10)$ times the mean heat flux at the core–mantle boundary.
@@ 1 Introduction
It is well-known that the Earth’s magnetic field is generated by thermochemical convection in the outer core. Compositional convection arises from the continual release of light elements that accompanies the progressive growth of the inner core. Thermal convection occurs primarily by secular cooling of the core as the planet cools in time. Mantle convection produces a laterally heterogeneous heat flux at the core–mantle boundary (CMB), which changes in geological time (Olson et al., 2010). Since the convective turnover time in the lower mantle is significantly longer than in the outer core (Zhang & Gubbins, 1992; Holme et al., 2015), these lateral variations create a quasi-stationary boundary condition that shapes outer core convection.
@@
Several observations indicate lower-mantle influence on the geodynamo. The persistence of high-latitude flux patch concentrations in the paleomagnetic time average (Gubbins & Kelly, 1993; Carlut & Courtillot, 1998; Johnson et al., 2003), the preferred paths of virtual geomagnetic poles during polarity transitions (Laj et al., 1991; Love, 2000), and the low secular variation in the Pacific (Fisk, 1931; Doell & Cox, 1971) are examples. Paleomagnetic data also suggest that the frequency of geomagnetic reversals changes on the time scale of mantle convection (Jones, 1977; McFadden & Merrill, 1984; Larson & Olson, 1991; Lhuillier et al., 2013). In addition, recent laboratory experiments (Sahoo & Sreenivasan, 2020) indicate that a large lower-mantle heterogeneity gives rise to an east–west dichotomy in core convection, which in turn explains the relative instability of the high-latitude magnetic flux lobes in the Western hemisphere (Jackson et al., 2000).
@@
The temperature variations in the lowermost mantle are based on the shear wave velocity anomalies obtained from seismic tomography, under the assumption that shear wave velocity is determined by temperature and not by composition (Su et al., 1994; Masters et al., 1996). The first self-consistent dynamo models with inhomogeneous CMB heat flux (Glatzmaier et al., 1999) showed that the frequency of geomagnetic reversals depends on the heat flux pattern. Since then, several numerical dynamo models (Olson & Christensen, 2002; Christensen & Olson, 2003; Aubert et al., 2007; Davies et al., 2008; Sreenivasan & Gubbins, 2011; Olson & Amit, 2014; Mound & Davies, 2023) have used CMB heat flux variations that linearly correlate with seismic shear wave velocity anomalies. Additionally, various single harmonics have been used to understand their effect in isolation on convection and the magnetic field (Glatzmaier et al., 1999; Olson et al., 2010; Kutzner & Christensen, 2004; Takahashi et al., 2008; Coe & Glatzmaier, 2006; Courtillot & Olson, 2007; Stanley et al., 2008; Amit et al., 2011). Mantle convection models indicate that the dominant spherical harmonic in the mantle flow switches between non-axisymmetric $l=1$ and $l=2$ (Zhong et al., 2007; Yoshida, 2008; Olson et al., 2010), where $l$ is spherical harmonic degree. During supercontinent aggregation, the non-axisymmetric $l=1$ mantle flow tends to dominate, whereas $l=2$ becomes predominant after supercontinent breakup. Therefore, it makes sense to focus on the effect of non-axisymmetric degree 1 and 2 heat flux patterns at the outer boundary. Of these, equatorially symmetric conditions are not likely to trigger polarity transitions (e.g. Takahashi et al., 2008). That said, the equatorially anti-symmetric part of the lower-mantle heat flux can become significant for large lateral variations in boundary heat flux and thus influence the polarity of the dynamo. The present study analyses this problem in the rapidly rotating, strongly driven regime of a planetary dynamo.
@@
A heterogeneous heat flux pattern at the outer boundary induces a steady mean flow, termed the thermal wind, which substantially changes the basic state temperature distribution of a rapidly rotating fluid layer. A study of the initiation of convection in the presence of the thermal wind (Teed et al., 2010) suggests that dynamo action may be supported in conditions that are otherwise subadiabatic. The heterogeneity that supports a dynamo is likely equatorially symmetric whereas an anti-symmetric condition inhibits dynamo action (Sahoo et al., 2016). In a thermally driven dynamo, an equatorially anti-symmetric boundary heat flux induces a non-zero mean axial temperature gradient $\partial T_{0}/\partial z$ at the equator, which modifies the buoyancy profile of the homogeneous state such that the resultant buoyancy frequency $\omega_{A}$ matches the Alfvén wave frequency $\omega_{M}$. This state marks the suppression of the slow Magnetic-Archimedean-Coriolis (MAC), or magnetostrophic, waves in the dynamo, in turn resulting in the collapse of the axial dipole (Majumder et al., 2024). The complementarity of the vertical (radial) and horizontal (lateral) buoyancies can in principle produce polarity transitions with weakly driven thermal convection and a large heterogeneity at the boundary. For a fixed vertical buoyancy, polarity reversals may exist in a narrow range of horizontal buoyancy states that lie between the dipolar and multipolar regimes. As in dynamos with uniform boundary heat flux (Majumder et al., 2024), we anticipate that the dipole–multipole transition would be self-similar, in that the resultant Rayleigh number based on the characteristic energy-containing length scale bears the same linear relationship with the square of the peak magnetic field at the transition, regardless of the scale of energy injection.
@@
The primary source of buoyancy for outer core convection is compositional, contributing up to 80% of the total convective power (Lister & Buffett, 1995). The weaker source of buoyancy, derived from thermal convection, is influenced by heat flux heterogeneity at the CMB. Compositional convection, arising from freezing of the inner core, exhibits long-term stability. In contrast, thermal convection, affected by variations in the CMB heat flux distribution, varies in geological time. Thus, thermal convection determines whether the dynamo maintains a stable dipole or transitions into a reversing regime, given that the intensity of compositional convection remains relatively invariant. In two-component convection, a complementarity exists between the vertical buoyancies of composition and temperature and the horizontal buoyancy of temperature. Using this and the condition for vanishing slow MAC waves, an upper bound for the lower-mantle heterogeneity that would admit an axial dipole field may be obtained.
@@
In §2, we consider the evolution of an isolated density disturbance in an unstably stratified fluid subject to a uniform magnetic field, background rotation and a lateral temperature variation. Here, the gravity and magnetic field axes coincide in Cartesian geometry, which is referred to as the ‘equatorial radial configuration’ by Loper et al. (2003). This analysis quantifies the resultant buoyancy based on the vertical and horizontal buoyancies and shows their complementary role in the evolution of fast and slow MAC waves. The Cartesian magnetoconvection model with non-zero $\partial T_{0}/\partial z$ serves as the basis for the study on the effect of equatorially anti-symmetric, or symmetry-breaking, outer boundary heat flux on wave motions in rapidly rotating spherical dynamos, described in §3. In a dynamo where the nonlinear inertial forces are small, an equatorially anti-symmetric condition causes polarity transitions not by breaking the symmetry of the columnar vortices but rather by selectively suppressing the slow MAC waves that already exist in the dipolar dynamo. This study also considers composite boundary heat flux patterns obtained by adding equatorially symmetric and antisymmetric patterns in a known proportion in order to extend the analysis to lower-mantle heat flux anomalies. §4 analyses two-component linear magnetoconvection and suggests an upper bound for the horizontal buoyancy that would admit a dipolar dynamo for a given thermal power ratio. In §5, we discuss the implications of our results for Earth and propose future work.
@@ 2.1 Problem set-up and governing equations
Figure 1: The initial state of a density perturbation $\rho^{\prime}$ that evolves in an unstably stratified fluid subject to a uniform magnetic field $\bm{B}=B\hat{\bm{e}}_{y}$, background rotation $\bm{\varOmega}=\varOmega\hat{\bm{e}}_{z}$, and gravity $\bm{g}=-g\hat{\bm{e}}_{y}$ in Cartesian coordinates $(x,y,z)$. The lateral variation in temperature produces a mean flow $\bm{u}_{0}=u_{c}z\,\hat{\bm{e}}_{x}$.
@@
$T=T_{0}+\Theta,\quad\bm{U}=\bm{u}_{0}+\bm{u},\quad\bm{B}=\bm{B}_{0}+\bm{b},$ (6a–c)
@@
where $\bm{j}$ is the electric current density, where $\nu$ is the kinematic viscosity, $\kappa$ is the thermal diffusivity, $\eta$ is the magnetic diffusivity and $p^{*}=p-(\rho/2)|{\bm{\Omega}}\times{\bm{x}}|^{2}$ is a modified pressure.
@@ 2.2 Solutions for the perturbation velocity field
where $k^{2}=k_{y}^{2}+k_{z}^{2}$ and the substitution $\hat{u}_{y}=-\hat{u}_{z}\,k_{z}/k_{y}$ is made from the continuity equation (10). Seeking a plane wave solution of the form $\hat{u}_{z}\sim\mbox{e}^{\mathrm{i}\lambda t}$, we obtain,
@@
[…]
represent the squares of the frequencies of linear inertial waves, vertical and horizontal parts of the buoyancy, Alfvén waves, magnetic diffusion, thermal diffusion and viscous diffusion, respectively. Here, $V_{M}=B_{0}/\sqrt{\rho\mu}$ is the Alfvén wave velocity. In line with a recent study (Sreenivasan & Maurya, 2021) where both viscous and thermal diffusion are much smaller than magnetic diffusion, the characteristic equation has the following form in the limit of $\omega_{\kappa},\omega_{\nu}\to 0$:
@@
where the coefficients $D_{m}$ are evaluated from the initial conditions of $\hat{u}_{z}$ and its derivatives (§$\S$2.3 below). Of the five terms in the expansion on the right-hand side of equation (20), two terms represent oppositely travelling fast MAC waves, two other terms represent oppositely travelling slow MAC waves, and the fifth term represents the overall growth of the velocity perturbation.

### v2

@@ Abstract
This study investigates, analytically and numerically, the role of a non-axisymmetric equatorially anti-symmetric lateral variation in heat flux at the outer boundary in polarity transitions in rapidly rotating dynamos. In an unstably stratified fluid, the frequencies of vertical and horizontal (lateral) buoyancy complement each other such that a polarity transition is induced by the boundary anti-symmetry through the suppression of the slow magnetic-Archimedean-Coriolis (MAC) waves at relatively small vertical buoyant forcing. A dipole-dominated dynamo in the low-inertia limit transitions to polarity-reversing and multipolar states in succession as the relative intensity of horizontal buoyancy is progressively increased for a fixed vertical buoyancy. In the same parameter space, an equatorially symmetric heat flux variation does not induce a polarity transition. A composite boundary heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a horizontal buoyancy of the same order as that for the transition induced by a purely anti-symmetric variation. This makes the analysis relevant to Earth’s core, which convects in response to a composite heat flux variation in the lowermost mantle. A heterogeneity with a dominant equatorially symmetric component does not favour polarity transitions, likely producing long periods in Earth’s history without reversals. The fact that compositional buoyancy is much stronger than thermal buoyancy, together with the known order of magnitude of the peak field intensity in the inertia-free limit, indicates a large lower-mantle heat flux heterogeneity of $O(10)$ times the mean superadiabatic heat flux at the core–mantle boundary.
@@ 1 Introduction
It is well-known that the Earth’s magnetic field is generated by thermochemical convection in the outer core. Compositional convection arises from the continual release of light elements that accompanies the progressive growth of the inner core. Thermal convection occurs primarily by secular cooling of the core as the planet cools in time. Mantle convection produces a laterally heterogeneous heat flux at the core–mantle boundary (CMB), which changes in geological time (Olson et al., 2010). Since the convective turnover time in the lower mantle is significantly longer than in the outer core (Zhang and Gubbins, 1992; Holme et al., 2015), these lateral variations create a quasi-stationary boundary condition that shapes outer core convection.
@@
Based on several observations, it has been argued that the lower mantle may influence the geodynamo. The persistence of high-latitude flux patch concentrations in the paleomagnetic time average (Gubbins and Kelly, 1993; Carlut and Courtillot, 1998; Johnson et al., 2003), the preferred paths of virtual geomagnetic poles during polarity transitions (Laj et al., 1991; Love, 2000), and the low secular variation in the Pacific (Fisk, 1931; Doell and Cox, 1971) are examples. Paleomagnetic data also suggest that the frequency of geomagnetic reversals changes on the time scale of mantle convection (Jones, 1977; McFadden and Merrill, 1984; Larson and Olson, 1991; Lhuillier et al., 2013). In addition, recent laboratory experiments (Sahoo and Sreenivasan, 2020) indicate that a large lower-mantle heterogeneity gives rise to an east–west dichotomy in core convection, which in turn explains the relative instability of the high-latitude magnetic flux lobes in the Western hemisphere (Jackson et al., 2000).
@@
The first self-consistent dynamo models with inhomogeneous CMB heat flux (Glatzmaier et al., 1999) showed that the frequency of geomagnetic reversals depends on the heat flux pattern. This CMB heat flux variation was estimated from the shear wave velocity anomalies obtained from seismic tomography, under the assumption that shear wave velocity is determined by temperature and not by composition (Su et al., 1994; Masters et al., 1996). However, compositional heterogeneity can also cause significant seismic wave velocity variations in the lower mantle (Koelemeijer et al., 2012; Gülcher et al., 2021). Several numerical dynamo models (Olson and Christensen, 2002; Christensen and Olson, 2003; Aubert et al., 2007; Sreenivasan and Gubbins, 2011; Olson and Amit, 2014; Mound and Davies, 2023) have used CMB heat flux variations that linearly correlate with seismic shear wave velocity anomalies. Additionally, various single harmonics have been used to understand their effect in isolation on convection and the magnetic field (Kutzner and Christensen, 2004; Takahashi et al., 2008; Coe and Glatzmaier, 2006; Courtillot and Olson, 2007; Stanley et al., 2008; Amit et al., 2011). It has been shown that both an enhanced equatorial heat flux (Glatzmaier et al., 1999; Olson et al., 2010) and a North–South hemispherical heat flux variation (Frasson et al., 2025) can produce reversals in numerical dynamos. Mantle convection models indicate that the dominant spherical harmonic in the mantle flow switches between non-axisymmetric $l=1$ and $l=2$ (Zhong et al., 2007; Yoshida, 2008; Olson et al., 2010), where $l$ is spherical harmonic degree. During supercontinent aggregation, the non-axisymmetric $l=1$ mantle flow tends to dominate, whereas $l=2$ becomes predominant after supercontinent breakup. Therefore, it makes sense to focus on the effect of non-axisymmetric degree 1 and 2 heat flux patterns at the outer boundary. Of these, equatorially anti-symmetric conditions are likely to trigger polarity transitions (e.g. Takahashi et al., 2008). The equatorially anti-symmetric part of the lower-mantle heat flux can become significant for large lateral variations in boundary heat flux and thus influence the polarity of the dynamo. The present study analyses this problem in the rapidly rotating, strongly driven regime of a planetary dynamo.
@@
A heterogeneous heat flux pattern at the outer boundary induces a steady mean flow, termed the thermal wind. A study of the onset of convection in a plane layer in the presence of a thermal wind (Teed et al., 2010) suggests that dynamo action might be possible even in subadiabatic conditions. Near convective onset, an equatorially symmetric boundary heterogeneity in heat flux likely favours dynamo action at large magnetic Prandtl number whereas an equatorially anti-symmetric heterogeneity can induce polarity reversals (Sahoo et al., 2016). In a thermally driven dynamo, an equatorially anti-symmetric heat flux heterogeneity induces a mean axial temperature gradient at the equator, which modifies the buoyancy profile of the homogeneous state, and hence the buoyancy frequency. For a sufficiently large anti-symmetric heterogeneity, the resultant buoyancy frequency $\omega_{A}$ matches the Alfvén wave frequency $\omega_{M}$, causing the suppression of the slow Magnetic–Archimedean–Coriolis (MAC), or magnetostrophic, waves in the dynamo. In a recent study (Majumder et al., 2024), variations in the mean outer boundary heat flux were interpreted as changes in the Rayleigh number, and the condition $|\omega_{A}|\approx|\omega_{M}|$ was shown to be associated with the loss of kinetic helicity from these waves, in turn leading to the collapse of the axial dipole magnetic field.
In an unstably stratified fluid, convection occurs when a superadiabatic gradient is present in the basic state. Within this medium, isolated density perturbations give rise to fast and slow MAC waves whose frequencies are obtainable in the Boussinesq limit (Braginsky, 1967; Busse et al., 2007; Majumder et al., 2024). In a rapidly rotating fluid where the magnitude of the inertial wave frequency $|\omega_{C}|\gg|\omega_{M}|,\,|\omega_{A}|$, the real part of the slow wave frequency is approximated by (e.g. Majumder et al., 2024),
$\omega_{s}\approx\dfrac{\omega_{M}^{2}}{\omega_{C}}\left(1+\dfrac{\omega_{A}^{2}}{\omega_{M}^{2}}\right)^{1/2}.$ (1)
The spontaneous generation of slow MAC waves in a convection-driven dynamo that evolves from a small seed magnetic field (figure S1, Supplementary material) indicates that these waves exist in an unstably stratified fluid where $\omega_{A}^{2}<0$, $\omega_{M}^{2}>0$ and $|\omega_{M}|\geq|\omega_{A}|$ in (1). As the dynamo passes through a chaotic multipolar state, the slow waves are excited by localized balances between the Lorentz, buoyancy and Coriolis forces, where the Lorentz force is made up of the non-dipolar field (figure S2, Supplementary Material). The helicity of the slow waves, which is at least as high as that of the fast waves, is essential for the formation of the dipole field from this multipolar state (Varma and Sreenivasan, 2022). The hydrodynamic dynamo, where the Lorentz force is absent, does not produce the dipole from a seed field (Sreenivasan and Kar, 2018) since the slow MAC waves are not generated in the first place. As the buoyant forcing is progressively increased in the dipolar regime, $|\omega_{M}|$ attains its highest value, upon which a state where $|\omega_{A}|\approx|\omega_{M}|$ ensues. Here, (1) predicts the disappearance of the slow MAC waves, which brings about the collapse of the axial dipole field (Majumder et al., 2024).
The present study builds on previous work, with a focus on examining how heterogeneity in the outer boundary heat flux causes the dipole–multipole transition. Since polarity transitions occur when $|\omega_{A}|\approx|\omega_{M}|$, and the resultant buoyancy frequency $\omega_{A}$ is made up of both vertical (radial) and horizontal (lateral) parts, it is in principle possible for a weak vertical buoyancy to be compensated by a strong horizontal buoyancy in order to match $|\omega_{M}|$. This complementarity between vertical and horizontal buoyancy suggests that polarity transitions can arise even under weakly driven thermal convection, provided there is a sufficiently large heat flux heterogeneity. For a fixed vertical buoyancy, polarity reversals may exist in a narrow range of horizontal buoyancy states that lie between the dipolar and multipolar regimes. As in dynamos with uniform boundary heat flux (Majumder et al., 2024), we anticipate that the dipole–multipole transition would be self-similar, in that the resultant Rayleigh number based on the characteristic energy-containing length scale bears the same linear relationship with the square of the peak magnetic field at the transition, regardless of the scale of energy injection.
@@
The formation and growth of the inner core make the primary source of buoyancy for outer core convection compositional, which is thought to contribute up to 80% of the total convective power (Lister and Buffett, 1995). The weaker source of buoyancy, derived from thermal convection, is influenced by heat flux heterogeneity at the CMB. In two-component convection, the vertical buoyancies of composition and temperature, together with the horizontal buoyancy from the lateral heat flux variation, make up the resultant buoyant forcing. Using this complementarity of buoyancies and the condition for vanishing slow MAC waves, an upper bound for the lower-mantle heterogeneity that would admit an axial dipole field may be obtained.
@@
In §2, we consider the evolution of an isolated density disturbance in an unstably stratified fluid subject to a uniform magnetic field, background rotation and a lateral temperature variation. Here, the gravity and magnetic field axes coincide in Cartesian geometry, which is referred to as the ‘equatorial radial configuration’ by Loper et al. (2003). This analysis quantifies the resultant buoyancy based on the vertical and horizontal buoyancies and shows their complementary role in the evolution of fast and slow MAC waves. The Cartesian magnetoconvection model with non-zero mean axial temperature gradient serves as the basis for the study on the effect of equatorially anti-symmetric, or symmetry-breaking, outer boundary heat flux on wave motions in rapidly rotating spherical dynamos, described in §3. In a dynamo where the nonlinear inertial forces are small, an equatorially anti-symmetric condition causes polarity transitions not by breaking the symmetry of the columnar vortices but rather by selectively suppressing the slow MAC waves that already exist in the dipolar dynamo. This study also considers composite boundary heat flux patterns obtained by adding equatorially symmetric and anti-symmetric patterns in a known proportion in order to extend the analysis to lower-mantle heat flux anomalies. §4 analyses two-component linear magnetoconvection and suggests an upper bound for the horizontal buoyancy that would admit a dipolar dynamo for a given thermal power ratio. In §5, we discuss the implications of our results for Earth and propose future work.
@@ 2.1 Problem set-up and governing equations
Figure 1: The initial state of a density perturbation $\rho^{\prime}$ that evolves in an unstably stratified fluid subject to a uniform magnetic field $\bm{B}_{0}=B_{0}\hat{\bm{e}}_{y}$, background rotation $\bm{\varOmega}=\varOmega\hat{\bm{e}}_{z}$, and gravity $\bm{g}=-g\hat{\bm{e}}_{y}$ in Cartesian coordinates $(x,y,z)$. The lateral variation in temperature produces a mean flow $\bm{u}_{0}=u_{c}z\,\hat{\bm{e}}_{x}$.
@@
$T=T_{0}+\Theta,\quad\bm{U}=\bm{u}_{0}+\bm{u},\quad\bm{B}=\bm{B}_{0}+\bm{b},$ (7a–c)
@@
where $\bm{j}$ is the electric current density, $\nu$ is the kinematic viscosity, $\kappa$ is the thermal diffusivity, $\eta$ is the magnetic diffusivity and $p^{*}=p-(\rho/2)|{\bm{\Omega}}\times{\bm{x}}|^{2}$ is a modified pressure.
@@ 2.2 Solutions for the perturbation velocity field
where $k^{2}=k_{y}^{2}+k_{z}^{2}$ and the following substitution is made from the continuity equation (11):
$\hat{u}_{y}=-\frac{k_{z}}{k_{y}}\hat{u}_{z}.$ (18)
Seeking a plane wave solution of the form $\hat{u}_{z}\sim\mbox{e}^{\mathrm{i}\lambda t}$, we obtain,
@@
$\begin{aligned} &\omega_{C}^{2}=4\varOmega^{2}k_{z}^{2}/k^{2},\quad\omega_{A,V}^{2}=g\alpha\beta_{y}k_{z}^{2}/k^{2},\quad\omega_{A,H}^{2}=g\alpha\beta_{z}k_{y}k_{z}/k^{2},\\ &\omega_{M}^{2}=B_{0}^{2}\,k_{y}^{2}/\rho\mu=V_{M}^{2}k_{y}^{2},\quad\omega_{\eta}^{2}=\eta^{2}k^{4},\quad\omega_{\kappa}^{2}=\kappa^{2}k^{4},\quad\omega_{\nu}^{2}=\nu^{2}k^{4},\end{aligned}$ (20a–g)
represent the squares of the frequencies of linear inertial waves, vertical and horizontal parts of the buoyancy, Alfvén waves, magnetic diffusion, thermal diffusion and viscous diffusion, respectively. Here, $V_{M}=B_{0}/\sqrt{\rho\mu}$ is the Alfvén wave velocity. In line with a recent study (Sreenivasan and Maurya, 2021) where both viscous and thermal diffusion are much smaller than magnetic diffusion, the characteristic equation has the following form in the limit of $\omega_{\kappa},\omega_{\nu}\to 0$:
@@
where the coefficients $D_{m}$ are evaluated from the initial conditions of $\hat{u}_{z}$ and its derivatives ($\S$2.3 below). Of the five terms in the expansion on the right-hand side of equation (22), two terms represent oppositely travelling fast MAC waves, two other terms represent oppositely travelling slow MAC waves, and the fifth term represents the overall growth of the velocity perturbation.